# Pathology-CoT: Learning Visual Chain-of-Thought Agent from Expert Whole Slide Image Diagnosis Behavior

### v1

@@ A.4 Prompts for VLM Analysis
Prompt for Final Summary and Diagnosis Generation Please provide a comprehensive final pathological impression and diagnosis based on all the above analyses. Consider: 1. The overall tissue architecture and morphology 2. Findings from each specific region 3. Any patterns or correlations between regions 4. Your final diagnostic impression, please do not consider suspicious region as positive Format your response EXACTLY as follows: <final_impression>Your comprehensive final diagnostic impression</final_impression> <recommendations>Any additional recommendations</recommendations> <diagnostic_info> PT_or_LN: "PT" if this is a primary tumor section, "LN" if this is a lymph node section t_stage: [1-4] if primary tumor, 0 if lymph node lymph_node_positive: true/false positive_regions: if lymph_node_positive is true, give a [1,2,3] like list. if false just say [] suspicious_regions: [] </diagnostic_info>
A.5 Method for Structuring Semi-Automated Thinking Labels
@@
The classification taxonomy was designed to be comprehensive for colorectal cancer lymph node metastasis. For the low-magnification (box-level) view, classes are grouped into three categories: (1) Tumor-related features (e.g., "Gland formation", "Necrosis"), (2) Normal Lymph Node Structures (e.g., "Germinal center", "Sinus"), and (3) Reactive or Pathological Changes (e.g., "Fibrosis", "Hemorrhage"). This allows for capturing the overall tissue architecture. For the high-magnification (40x-level) view, classes are also grouped into three categories: (1) Tumor Cell-related features (e.g., "Tumor cell", "Mitotic figure"), (2) Normal Background Cells and Structures (e.g., "Lymphocyte", "Extracellular matrix"), and (3) Pathological Processes (e.g., "Apoptotic body", "Dead cell"). This captures fine-grained cytological details.
@@
Prompt for Multi-Label Classification of Thinking Labels Based on the conversation text, classify what is observed in this ROI for both low-magnification (box) and high-magnification (40x) views. This is a MULTI-LABEL classification - multiple classes can be present simultaneously. Box classes (low-magnification): Tumor deposit, Gland formation, Tumor stroma, Necrosis, Germinal center, Lymphoid follicle, Medullary cord, Sinus, Paracortex, Sinus histiocytosis, Fibrosis, Congestion, Hemorrhage, Fatty replacement, other 40x classes (high-magnification): Tumor cell, Mitotic figure, Atypical glandular cell, Signet ring cell, Lymphocyte, Plasma cell, Macrophage, Endothelial cell, Fibroblast, Erythrocyte, Fat cell, Extracellular matrix, Apoptotic body, Dead cell, Inflammatory cell, other Return ONLY the two classification results in this format: box_classifications|40x_classifications Where each classification is a comma-separated list of applicable classes. Conversation text: text Instructions: - Choose ALL applicable classes for each view from the provided lists - Multiple classes can be present simultaneously (e.g., "Gland formation,Fibrosis") - If none match well, use "other" - Return only the two classification lists separated by |
A.6 Specification of Evaluated Vision-Language Models

### v2

@@ A.2 Prompts for VLM Analysis
Prompt for Final Summary and Diagnosis Generation Please provide a comprehensive final pathological impression and diagnosis based on all the above analyses. Consider: 1. The overall tissue architecture and morphology 2. Findings from each specific region 3. Any patterns or correlations between regions 4. Your final diagnostic impression, please do not consider suspicious region as positive Format your response EXACTLY as follows: <final_impression>Your comprehensive final diagnostic impression</final_impression> <recommendations>Any additional recommendations</recommendations> <diagnostic_info> PT_or_LN: “PT” if this is a primary tumor section, “LN” if this is a lymph node section t_stage: [1-4] if primary tumor, 0 if lymph node lymph_node_positive: true/false positive_regions: if lymph_node_positive is true, give a [1,2,3] like list. if false just say [] suspicious_regions: [] </diagnostic_info>
A.3 Method for Structuring Semi-Automated Thinking Labels
@@
The classification taxonomy was designed to be comprehensive for colorectal cancer lymph node metastasis. For the low-magnification (box-level) view, classes are grouped into three categories: (1) Tumor-related features (e.g., “Gland formation”, “Necrosis”), (2) Normal Lymph Node Structures (e.g., “Germinal center”, “Sinus”), and (3) Reactive or Pathological Changes (e.g., “Fibrosis”, “Hemorrhage”). This allows for capturing the overall tissue architecture. For the high-magnification (40x-level) view, classes are also grouped into three categories: (1) Tumor Cell-related features (e.g., “Tumor cell”, “Mitotic figure”), (2) Normal Background Cells and Structures (e.g., “Lymphocyte”, “Extracellular matrix”), and (3) Pathological Processes (e.g., “Apoptotic body”, “Dead cell”). This captures fine-grained cytological details.
@@
Prompt for Multi-Label Classification of Thinking Labels Based on the conversation text, classify what is observed in this ROI for both low-magnification (box) and high-magnification (40x) views. This is a MULTI-LABEL classification - multiple classes can be present simultaneously. Box classes (low-magnification): Tumor deposit, Gland formation, Tumor stroma, Necrosis, Germinal center, Lymphoid follicle, Medullary cord, Sinus, Paracortex, Sinus histiocytosis, Fibrosis, Congestion, Hemorrhage, Fatty replacement, other 40x classes (high-magnification): Tumor cell, Mitotic figure, Atypical glandular cell, Signet ring cell, Lymphocyte, Plasma cell, Macrophage, Endothelial cell, Fibroblast, Erythrocyte, Fat cell, Extracellular matrix, Apoptotic body, Dead cell, Inflammatory cell, other Return ONLY the two classification results in this format: box_classifications|40x_classifications Where each classification is a comma-separated list of applicable classes. Conversation text: text Instructions: - Choose ALL applicable classes for each view from the provided lists - Multiple classes can be present simultaneously (e.g., “Gland formation,Fibrosis”) - If none match well, use “other” - Return only the two classification lists separated by |
A.4 Specification of Evaluated Vision-Language Models